A small object of mass 15 g carries a charge 3400 C and is suspended by a thread between the vertical plates of a parallel-plate capacitor. The plate separation is 3.55 cm. If the thread makes an angle of 10 degrees with the vertical, what is the potential difference between the plates?
1
Expert's answer
2020-11-05T10:36:46-0500

For Vertical Equilibrium-

Tcos(10°) = mg

0.9848T = (0.015)(9.8)

Tension in thread (T) = 0.149 N

For Horizontal Equilibrium-

FQ = Tsin(10°)

qE = (0.1737)(0.149)

(3400)("\frac{V}{d}") = 0.02587

"V = \frac{0.02587 \times 0.0355 }{3400}"

"V = 2.7 \times 10^{-7} V"